OpenStudy (anonymous):

The quadratic formula. Solve. t2 + 10 = 6t = 1t^2 + 10 - 6t= 0 Not sure where to go from here. any help?

OpenStudy (anonymous):

\[Quadratic Formula: (-b \pm \sqrt{b ^{2}-4ac})/2a\] Then just plug in all the numbers.

OpenStudy (chaise):

I helped you with this one like 8 hours ago. Check your notifications.

OpenStudy (anonymous):

Not sure what happen but I don't see any notifications.

OpenStudy (chaise):

jbmims25 replied to The quadratic formula. Solve. 19. t2 + 10 = 6t 8 hours ago Go look for it :)

OpenStudy (anonymous):

Make sure the equation is in the form ax^2+bx+c=0 before using the formula, which you did except you could re-write it as: t^2-6t+10=0, just so you can clearly see what a, b, & c are. a=1, b=-6 (include negative sign), c=10 Then just plug and chug in the quadratic formula :) Things I would look out for are your + and - signs... whether or not you are subtracting a negative (will make the number a positive), squaring a negative number (also makes it positive), etc. Also it looks like you're going to have an imaginary number in your answer (square root of -1 = i).

OpenStudy (anonymous):

\[-(-6) \pm \sqrt{(-6)^{2}-4(1)(10)} \] /2(1)

OpenStudy (anonymous):

\[6\pm \sqrt{36-40}\] /2 = \[6\pm \sqrt{-4}\] /2

OpenStudy (anonymous):

The square root of (-4) can be viewed as the square root of (-1)*(4), so since you can take the square root of products you'll have 2i as the answer to the square root: \[6\pm2\iota \] /2 If you really want to reduce this, you can divide both of the top terms by 2 and get: \[3\pm \iota\] (this is the same as above because it is a sum of two fractions: \[(6/2) \pm (2i/2) = (6\pm2i)/2\] Don't know how in depth you needed but hope it helps :)